OpenStudy (richhomiehelp):

What is the approximate volume of a cylinder with a height of 2 ft and a radius of 6 ft? Use 3.14 to approximate pi. Express your answer in hundredths.

OpenStudy (unklerhaukus):

The are of a cylinder is equal to the area of its base, times its height. The base is a circle.

OpenStudy (unklerhaukus):

The area of a circle is π r^2, where r is the radius.

OpenStudy (richhomiehelp):

Im So Confused @UnkleRhaukus

OpenStudy (unklerhaukus):

Can you compute the area of the circle? A = π r^2 = 3.14 x ( 6 ft )^2 =

OpenStudy (richhomiehelp):

10.5

OpenStudy (unklerhaukus):

did you square the 6?

OpenStudy (richhomiehelp):

idk how

OpenStudy (unklerhaukus):

6^2 = 6 x 6

OpenStudy (richhomiehelp):

ooo yea

OpenStudy (unklerhaukus):

so you have A = π r^2 = 3.14 x ( 6 ft )^2 = 3.14 x 6 x 6 ft^2 =

OpenStudy (richhomiehelp):

113.04

OpenStudy (unklerhaukus):

great!, so the area of the circular base is 113.04 ft^2

OpenStudy (unklerhaukus):

Now we can find the volume of the cylinder, but multiplying by the height

OpenStudy (richhomiehelp):

Thanks

OpenStudy (richhomiehelp):

ok

OpenStudy (unklerhaukus):

What do you get?

OpenStudy (richhomiehelp):

idk

OpenStudy (richhomiehelp):

ooo

OpenStudy (unklerhaukus):

V = 113.04 ft^2 x 2 ft = 113.04 x 2 ft^3 = . . .

OpenStudy (richhomiehelp):

226.08

OpenStudy (unklerhaukus):

That's it.

OpenStudy (richhomiehelp):

THANKS!!

OpenStudy (richhomiehelp):

1 More

OpenStudy (unklerhaukus):

the units are ft^3.

OpenStudy (richhomiehelp):

Yes

OpenStudy (unklerhaukus):

ok, next one

OpenStudy (richhomiehelp):

The radius of a cylinder is 15 cm. The volume of the cylinder, found using 3.14 to approximate pi, is 19,782 cm3. What is the height of the cylinder? Express your answer as a whole number.

OpenStudy (unklerhaukus):

As before, for a cylinder the volume is V = Ah = 2 π r^2 h Solve V = 2π r^2 h for h.

OpenStudy (richhomiehelp):

???

OpenStudy (unklerhaukus):

find an equivalent equation that is like h = ........

OpenStudy (unklerhaukus):

V = 2π r^2 h divide both sides of the equation by 2 π r^2

OpenStudy (richhomiehelp):

???????????

OpenStudy (unklerhaukus):

V / 2π r^2 = 2π r^2 h / 2π r^2

OpenStudy (unklerhaukus):

simplify the righthand side

OpenStudy (richhomiehelp):

19782

OpenStudy (unklerhaukus):

cancel the terms common to both the numerator and denominator

OpenStudy (richhomiehelp):

706

OpenStudy (unklerhaukus):

V / 2π r^2 = h or switching it around h = V / ( 2π r^2 ) now plug in the values